OpenStudy (anonymous):

PLsss i need Help! lim ( (2^n + 3^n)/ (5n-1) ) ^ 1/n n goes infinty

OpenStudy (anonymous):

\[(\frac{2^n+3^n}{5n-1})^\frac{1}{n}\]

OpenStudy (anonymous):

\[A=\lim_{x \rightarrow \infty} \left\{ (2^{x}+3^{x})/(5x-1)\right\}^{1/x}\] \[\ln A=\lim_{x \rightarrow \infty}(1/x)*\ln \left\{ (2^{x}+3^{x})/(5x-1)\right\}\] \[=\lim_{x \rightarrow \infty}\ln \left\{ (2^{x}+3^{x})/(5x-1)\right\}/x\] this is \[\infty/\infty\] \[=\lim_{x \rightarrow \infty}\ln \left\{ (2^{x}+3^{x})/x-\ln(5x-1)\right\}/x\] use l'Hospital's \[(2^{x}\ln2+3^{x}\ln3)/(2^{x}+3^{x})-5/(5x-1)\] \[\lim_{x \rightarrow \infty}(2^{x}\ln2)/(2^{x}+3^{x})+\lim_{x \rightarrow \infty}(3^{x}\ln3)/(2^{x}+3^{x})-\lim_{x \rightarrow \infty}5/(5x-1)\] first term is zero, second term is ln3 and third term is zero so lnA=ln3 A=3

OpenStudy (anonymous):

thx a lot good solving